OpenStudy (calculator):

Electrochemistry problem. A solution of iron(III) chloride is used to dissolve unwanted copper from printed circuit boards. When a copper-coated printed circuit board is immersed in FeCl 3(aq), the solution turns pale blue. Suggest an equation for the reaction between copper and iron(III) chloride and use the Standard electrode potential (data) to calculate the E° value for the reaction.

OpenStudy (calculator):

They have given me the correct reaction which is \[2Fe^{3+}+2e^-\rightarrow 2Fe^{2+},~E^\theta=+0.77V \\ \\ Cu \rightarrow Cu^{2+}+2e^-,~E^\theta =+0.34V\] E=0.43V but my thought is why can't we use \[Fe^{3+}+e^- \rightarrow Fe^{2+},~E^\theta=-0.04V\] \[Cu \rightarrow Cu^++e^-,~E^\theta=+0.52V\] E=-0.56V

OpenStudy (calculator):

@mukushla @Hero

OpenStudy (calculator):

@Xishem @salini

OpenStudy (anonymous):

here for conversion of elemental copper to ionic state we need 2 electrons cu- does not give the correct requirement here the other eqn gets multipleid by 2 as we get 2 electrons from cu to cu2+

OpenStudy (anonymous):

 note the 2+ oxidation state